OpenStudy (anonymous):

The formula for the surface area of a rectangular prism is a=2(lw+lh+wh) . Solve the formula for w

Directrix (directrix):

A=2(lw+lh+wh) A = 2*lw + 2*lh +2*wh The object is to get the w terms together on one side of the equation and the other non-w-containing terms on the other side. 2*lw + 2*wh = A - 2*lh 2w( l + h ) = A - 2*lh w = (A - 2lh) /[2(l + h)] --> Check this and see what you think.